# Bootstrapping Autonomous Radars with Self-Supervised Learning

### v1

@@ Abstract
The perception of autonomous vehicles using radars has attracted increased research interest due its ability to operate in fog and bad weather. However, training radar models is hindered by the cost and difficulty of annotating large-scale radar data. To overcome this bottleneck, we propose a self-supervised learning framework to leverage the large amount of unlabeled radar data to pre-train radar-only embeddings for self-driving perception tasks. The proposed method combines radar-to-radar and radar-to-vision contrastive losses to learn a general representation from unlabeled radar heatmaps paired with their corresponding camera images. When used for downstream object detection, we demonstrate that the proposed self-supervision framework can improve the accuracy of state-of-the-art supervised baselines by in mAP.
@@ 1 Introduction
However, compared to de facto computer vision datasets like ImageNet, the volume of annotated open radar datasets remains very limited. This is because radar images are especially challenging for humans to interpret and thus annotate. Fig. 1 shows an example of a bird’s eye view (BEV) radar heatmap and the corresponding camera image. Unlike camera images, radar heatmaps appear as blobs with no sharp boundaries or well-defined shapes for the objects present in the scene. These blobs carry little to no contextual or perceptual information and, as such, are hard to interpret by humans. Furthermore, mmWave radar signals are highly specular; meaning, mmWave signals exhibit mirror-like reflections on cars [12]. As a result, not all reflections from the car propagate back to the radar receiver, and most of the car does not appear in the image. These effects compound making it difficult even for well-trained radar imaging experts to draw precise bounding boxes of objects [25]. As a result, only a tiny fraction of radar data is typically labeled (e.g. 10%) of the hundreds of thousands of raw radar frames in open radar datasets [39]. Hence, building accurate supervised radar object detection models is extremely difficult.
To address the challenge of annotating radar data, prior work leverages other sensing modalities like cameras and lidar to derive labels for radar heatmaps and use these labels as ground truth to train radar-based models [47, 59, 61, 57, 36, 35, 22]. However, different sensory modalities have different viewpoints and projection planes of the scene. For example, camera-based labels suffer from depth-unaware perspective projection onto the image plane, so they cannot provide accurate supervision along the depth axis in bird’s eye view (BEV) radar heatmaps. Errors in viewpoint alignment between the different sensory modalities also result in highly in accurate detection. Moreover, because radar and optical sensors (camera and lidar) operate on orthogonal portions of the electromagnetic spectrum, objects that are visible to optical sensing are not necessarily visible to radar and vice versa. Directly using lidar data to supervise the training of radar will force the radar model to focus too much on less prominent reflections in radar heatmaps, such as less-visible surfaces due to specularity. In contrast, certain materials, such as glass, are not visible to optical sensors but are visible to radars. Therefore, cross-modal supervision results in false positive (FP) and false negative (FN) detection [36]. Finally, as radar hardware continues to evolve, it requires us to keep labeling new datasets collected using new radar hardware, which is going to be very expensive in the long run.
@@
Applying self-supervised learning (SSL), which has been extensively studied in the NLP and CV communities, to the radar domain, is not trivial because state-of-the-art self-supervised learning methods are designed for camera images. They either design pretext prediction tasks for RGB images [21, 30], or leverage camera-specific attributes to design strong augmentations to enforce semantic invariance [18, 19]. RGB augmentation methods cannot be generalized to RF sensing data, including radar. For example, radar data are natively associated with polar coordinates and hence are not invariant to transformations like translation and resizing. Previous work [37] on sensing the human pose found that directly applying popular self-supervised learning frameworks like [18, 29, 64] to radar heatmaps results in “shortcuts” in the learnt representation rather than capturing meaningful radar information.
@@
Second, we introduce a novel augmentation technique RMM (Radar MIMO Mask) that is tailored for state-of-the-art automotive radars. RMM leverages the fact that these radars use MIMO which combines multiple transmitters and multiple receivers. We manipulate how we combine the raw signals coming from different transmitter/receiver pairs to generate new augmented radar heatmaps. This augmentation preserves the underlying geometric structure of the scene while promoting resilience to the radar noise induced by Doppler phase distortions [26].
@@
To the best of our knowledge, this is the first work on autonomous driving that uses self-supervised learning to take advantage of the vast amounts of unlabeled radar data and achieve 2D bounding box detection using radar only. Our findings may prove key in generating pre-trained models that avoid the need to annotate massive amounts of radar data and enable lifelong learning on new radar hardware and datasets. Finally, we will make our code publicly available once the paper is accepted.
@@ 2 Related Work
Self-supervised learning. SSL, in its contrastive and non-contrastive flavours, has by now become a staple of representation learning for computer vision tasks [18, 29, 24, 69, 46, 17, 13, 23]. At the core of vision SSL lies augmentation for synthetically generating positive views for enforcing semantic invariance. We build on two pioneering contrastive SSL methods for vision: SimCLR and MoCo [18, 29, 20]. SimCLR introduced the canonical contrastive architecture using in-batch negative sampling, which typically relies on a large batch size and associated memory. MoCo uses an efficient queue and momentum update, which decouples negative sampling from the batch size. Although we heavily draw on vision SSL, our work recasts recent advances within a new cross-modal learning objective for accurate vision-free bounding box estimation.
@@ 4.1 Distillation Setup
Let be a radar-vision data pair in dataset , where is a radar heatmap and is a corresponding RGB image. Encode the radar heatmap with a backbone net then project it with an MLP head , assuming some weight parametrisation , such that . Similarly encode the paired visual image such that , with being a pretrained and frozen vision backbone model. Knowledge is distilled from the pretrained vision backbone and into the radar model by means of local interactions at the radar branch, as well as global interactions with the vision branch as depicted in Fig. 2.
@@ 4.3 Cross-modal Radar-vision Learning
Similar to the radar-to-radar contrastive learning term in Eq. 1, we then compute the term , where the use of the prototype in radar-to-vision and vision-to-radar contrastive terms is denoted by . The in-batch cross-modal contrastive loss is then given by:
@@ 4.4.1 Repurposed Vision Augmentations
Considering that BEV radar heatmaps have formats similar to camera images, a subset of standard SSL vision augmentations is potentially applicable to radar heatmaps. However, due to the naturally different perspectives and coordinate system, most vision augmentations are not applicable or need to be modified first.
We conduct extensive experiments on different vision augmentations and their combinations (cf., Sec. 6). We find that horizontal flip, rotation, and center cropping [18] are also suitable for radar heatmaps. We note that for radar heatmaps naturally in polar coordinates, rotation and center cropping should be applied in the polar coordinates, as shown in Fig. 3(f) and (g) respectively.
@@ 4.4.2 Radar-specific augmentations
In addition to the adoptable subset of standard SSL vision augmentations, we introduce and experiment with a new domain-specific augmentation for radar learning we call Radio MIMO Mask (RMM). First, we briefly explain how the raw data is processed before RMM is applied.
RMM Implementation. While raw radar data is usually processed to be presented as in the form of range-azimuth heat maps, point clouds, or range-Doppler maps [51, 55, 39], Radical uses an intermediate 3-D vector to apply RMM. Specifically, consider a MIMO radar with transmitters and receivers. A range-azimuth heat map of size is generated when a 3-D complex array of size is summed over all the antenna pairs (i.e., the first dimension). Radical’s RMM is applied right before this summation happens. RMM is best presented as the composition of two operations, antenna dropout and random phase noise, which we explain in more detail below.
(1) Antenna Dropout. We leverage the reconfigurability of the virtual array emulated by the MIMO radar to design this radar-specific augmentation. We omit randomly a subset of virtual antenna elements from the subsequent signal aggregation. Mathematically, we can write
where is the augmented radar heatmap as a function of range and azimuth angle , and iterates over the set of antenna pairs, and are independent discrete random masks that nullifies the antenna pair with probability . This augmentation simulates scenarios with partial sensor failure or obstructions, which promotes learning from incomplete data and improves robustness. The probability of antenna dropout is a tunable hyper-parameter.
@@
where is the signal from the transmitter and receiver pair, is the augmented signal, and are i.i.d. phase shifts drawn from uniform distributions between and (in radians), where is a tunable hyper-parameter. This randomization mimics the phase variability introduced by environmental factors and relative motions between the radar and the scene, which is also referred to as Doppler-induced phase noise [39, 26]. Thus it enhances the training coverage of RF conditions likely to occur in the real-world. We note that larger corresponds to more aggressive movements and noise in the environment.
RMM. The final RMM augmentation is the (order-agnostic) composition of the two operations above. We found empirically that and for the Bernoulli and the phase shift range hyper-parameters leads to the best performance in our experiments (see sec. 6.2).
4.5 Down-stream Fine-tuning
@@ 4.6 Implementation Details
For the radar backbone, we use Radatron [39], which adopts an FPN-based architecture. The backbone has a two-stream architecture, which takes as inputs high- and low-resolution radar heatmaps. Though generated from the cascade MIMO capture of the same scene, the high- and low-res streams undergo different radar signal pre-processing. Specifically, each stream goes through a stem layer and then two ResNet stages, which are identical to the building blocks of ResNet50 [28]. Then the two streams are concatenated and fused in a convolutional layer. The resultant feature maps are further encoded via additional ResNet stages, and combined to create the features similar to [63]. We pre-train the backbone of the model (without the FPN and the linear regression heads) as the radar feature extractor.
For the vision branch, we use a pre-trained CLIP image encoder model [52], which we freeze in pre-training.
@@ 5.1 Dataset
We evaluate Radical on few publicly available datasets that support raw radar data format. This is because our domain-specific augmentations require raw radar format. We focus here on using Radatron dataset [39], with a total of 152K radar frames and 16K labeled frames. In addition to the requisite raw radar format, we find Radatron’s size (both labeled and unlabeled frames) beneficial in the characterisation we present herein. Out of the unlabeled set, we use K frames for self-supervised pretraining, K annotated frames for supervised fine-tuning, and K annotated frames for testing. The train and test splits are constructed from experiments conducted on different days throughout the data collection campaign. The raw radar frames are first converted to complex, 86-channel heatmaps. These heatmaps are then fed to the network for pre-processing and stochastic augmentation.
@@ 5.2 Experiments
We pre-train Radical as depicted in Fig. 2 and detailed in Secs. 4. We utilize unlabeled radar-vision frames from Radatron as described in Sec. 5.1. We specialize in the pre-trained radar embeddings for a downstream task relevant to self-driving. The task strives to detect rotated 2D bounding boxes in the bird’s eye view (BEV) from radar heatmaps.
During pre-training, we use a batch size of 128, learning rate of 0.05 and cosine learning rate scheduling with an SGD optimizer with momentum 0.9 and weight decay 0.0001. During fine-tuning, we adopt the same training setting as Radatron, using a batch size of 8, an SGD optimizer with learning rate of 0.01 and 25K iterations with learning rate drop at 15K and 20K iterations. We increase the weight decay to 0.001 in order to avoid overfitting problems, boosting the baseline performance.
If not stated otherwise, our results are obtained with pre-training the backbone using 63K unlabeled frames, and fine-tuning the downstream model on 13K labeled frames. Results are averaged over 6 runs.
@@ 5.3 Baselines
(2) Intra-modal CL. We disable vision from contributing to the composite contrastive loss by setting in Eq. 4, which results in intra-modal, radar-only CL. For this, we use the vision-based augmentations of vertical flipping and center cropping.
@@ 6 Results
This section presents a comprehensive analysis of Radical’s performance against baselines and examines the impact of various augmentations on model performance. Evaluation Metrics. Following previous radar detection work [51, 39] we use Average Precision (AP) with IoU thresholds of , and to evaluate Radical’s detection performance. We also use the mean AP (mAP) of IoU values from 0.5 to 0.95 in 0.05 steps. We follow the COCO framework [38] for evaluating Radical.
@@
Fine-tuning. We pre-train Radical utilising our composite intra- and cross-modal CL configuration, along with the two baseline CL configurations from Sec. 6.1. We then fine-tune all pre-trained backbones along with their bounding box estimation heads. We compare these pre-trained variants to the implementation of [39] which uses random initialization. Table 1 shows the quantitative results using three metrics: mAP, AP, and AP. The mean and standard deviation of these results are obtained from 6 different runs of supervised training, while keeping the pre-trained weights the same. As we can seen, Radical’s composite intra- and cross-modal configuration performs most favourably, and outperforms random initialization by 5.8% in mAP. This demonstrates the efficacy of Radical’s pre-training on this highly relevant downstream task. Radical also outperforms intra-modal CL and cross-modal CL baselines by 2.9% and 2.4% respectively. Despite good gains over random initialisation (approx. 3% each), the two single-modal CL baselines are unable to approach the performance of Radical’s composite CL loss.
| Labeled data fraction | |||||
| Radatron [39] | |||||
| Radical (ours) | |||||
@@
Freezing. We fix the pre-trained parameters in order assess and compare the quality of the learnt features across our contrastive configurations. To this end, we similarly to above train task-specific heads for our downstream bounding box estimation task. We randomly initialize Radatron’s backbone and then similarly freeze it. The averages and standard deviations are listed in Table 2. We observe that Radical outperforms all baselines on all metrics. Random initialisation performs poorly compared to the pre-trained variants. This highlights the inadequacy of the task-specific head to perform accurate bounding box estimation without quality featurisation underneath. We also observe that the gap between Radical and the two CL baselines widens compared to Table 1. Without fine-tuning to compensate, this further underscores the efficacy of Radical compared to single-modal CL variants. We also observe a slight performance advantage to cross-modal over the intra-modal CL variants. This could point to the importance of visual priors in the training of quality radar embeddings.
Label Efficiency. We investigate the impact of the number of available labeled data on performance under a fine-tuning protocol. We compareRadical to the random initialization of Radatron. For Radical, we use all unlabeled data for pretraining. Table 3 shows the mAP after full fine-tuning as a function of the fraction of labeled data used. We see consistent improvements using Radical pre-training over the supervised baseline across all label density regimes.

### v3

# Bootstrapping Autonomous Driving Radars with Self-Supervised Learning

@@ Abstract
The perception of autonomous vehicles using radars has attracted increased research interest due its ability to operate in fog and bad weather. However, training radar models is hindered by the cost and difficulty of annotating large-scale radar data. To overcome this bottleneck, we propose a self-supervised learning framework to leverage the large amount of unlabeled radar data to pre-train radar-only embeddings for self-driving perception tasks. The proposed method combines radar-to-radar and radar-to-vision contrastive losses to learn a general representation from unlabeled radar heatmaps paired with their corresponding camera images. When used for downstream object detection, we demonstrate that the proposed self-supervision framework can improve the accuracy of state-of-the-art supervised baselines by in mAP. Code is available at https://github.com/yiduohao/Radical.
@@ 1 Introduction
However, compared to de facto computer vision datasets like ImageNet, the volume of annotated open radar datasets remains very limited. This is because radar images are especially challenging for humans to interpret and thus annotate. Figure 1 shows an example of bird’s eye view (BEV) radar heatmaps and the corresponding camera images. Unlike camera images, radar heatmaps appear as blobs with no sharp boundaries or well-defined shapes for the objects present in the scene. These blobs carry little to no contextual or perceptual information and, as such, are hard to interpret by humans. Furthermore, mmWave radar signals are highly specular; meaning, mmWave signals exhibit mirror-like reflections on cars [12]. As a result, not all reflections from the car propagate back to the radar receiver, and most of the car does not appear in the image. These effects compound making it difficult even for well-trained radar imaging experts to draw precise bounding boxes of objects [27]. As a result, only a tiny fraction of radar data is typically labeled (e.g. 10%) of the hundreds of thousands of raw radar frames in open radar datasets [43]. Hence, building accurate supervised radar object detection models is extremely difficult.
To address the challenge of annotating radar data, prior work leverages other sensing modalities like cameras and lidar to derive labels for radar heatmaps and use these labels as groundtruth to train radar-based models [51, 66, 68, 64, 39, 38, 23]. However, different sensory modalities have different viewpoints and projection planes of the scene. For example, camera-based labels suffer from depth-unaware perspective projection onto the image plane, so they cannot provide accurate supervision along the depth axis in BEV radar heatmaps. Errors in viewpoint alignment between the different sensory modalities also result in highly inaccurate detection. Moreover, because radar and optical sensors (camera and lidar) operate on orthogonal portions of the electromagnetic spectrum, objects that are visible to optical sensing are not necessarily visible to radar and vice versa. Directly using lidar data to supervise the training of radar will force the radar model to focus too much on less prominent reflections in radar heatmaps, such as less-visible surfaces due to specularity. In contrast, certain materials, such as glass, are not visible to optical sensors but are visible to radars. Therefore, cross-modal supervision results in false positive and false negative detection [39]. Finally, as radar hardware continues to evolve, it requires us to keep labeling new datasets collected using new radar hardware, which is going to be very expensive in the long run.
@@
Applying self-supervised learning (SSL), which has been extensively studied in the NLP and CV communities, to the radar domain, is nontrivial because state-of-the-art self-supervised learning methods are designed for camera images. They either design pretext prediction tasks for RGB images [22, 32], or leverage camera-specific attributes to design strong augmentations to enforce semantic invariance [19, 20, 55]. RGB augmentation methods cannot be generalized to RF sensing data, including radar. For example, radar data are natively associated with polar coordinates and hence are not invariant to transformations like translation and resizing. Previous work [40] on sensing the human pose found that directly applying popular SSL frameworks like [19, 31, 71] to radar heatmaps results in “shortcuts” in the learnt representation rather than capturing meaningful radar information.
@@
Second, we introduce a novel augmentation technique RMM (Radar MIMO Mask) that is tailored for state-of-the-art automotive radars. RMM leverages the fact that these radars use MIMO which combines multiple transmitters and multiple receivers. We manipulate how we combine the raw signals coming from different transmitter/receiver pairs to generate new augmented radar heatmaps. This augmentation preserves the underlying geometric structure of the scene while mimicing the radar noise induced by Doppler phase distortions [28].
@@
To the best of our knowledge, this is the first work on autonomous driving that uses self-supervised learning to take advantage of the vast amounts of unlabeled radar data and achieve 2D bounding box detection using radar only. Our findings may prove key in generating pre-trained models that avoid the need to annotate massive amounts of radar data and enable lifelong learning on new radar hardware and datasets.
@@ 2 Related Work
Self-supervised learning. SSL, in its contrastive and non-contrastive flavours, has by now become a staple of representation learning for computer vision tasks [19, 31, 26, 76, 50, 18, 13, 25, 42]. At the core of vision SSL lies augmentation for synthetically generating positive views for enforcing semantic invariance. We build on two pioneering contrastive SSL methods for vision: SimCLR and MoCo [19, 31, 21]. SimCLR introduced the canonical contrastive architecture using in-batch negative sampling, which typically relies on a large batch size and associated memory. MoCo uses an efficient queue and momentum update, which decouples negative sampling from the batch size. Although we heavily draw on vision SSL, our work recasts recent advances within a new cross-modal learning objective for accurate vision-free bounding box estimation.
@@ 4.1 Distillation setup
Let be a radar-vision data pair in dataset , where is a radar heatmap with range bins and azimuth bins, and is a corresponding RGB image. Encode the radar heatmap with a backbone net then project it with an MLP head , assuming some weight parametrisation , such that . Similarly encode the paired visual image such that , with being a pretrained and frozen vision backbone model. Knowledge is distilled from the pretrained vision backbone and into the radar model by means of local interactions at the radar branch, as well as global interactions with the vision branch as depicted in Fig. 2.
@@ 4.3 Cross-modal radar-vision learning
Similar to the radar-to-radar contrastive learning term in Eq. 1, we then compute the term , where the use of the prototype in radar-to-vision contrastive term is denoted by . The in-batch cross-modal contrastive loss is then given by
@@ 4.4.1 Repurposed vision augmentations
Considering that BEV radar heatmaps have formats similar to camera images, a subset of standard SSL vision augmentations is potentially applicable to radar heatmaps. However, due to the different perspectives and coordinate system, most vision augmentations are not applicable or need to be carefully modified.
We conduct extensive experiments on different vision augmentations and their combinations (cf., Sec. 6). We find that horizontal flip, rotation, and center cropping [19] are also suitable for radar heatmaps. We note that for radar heatmaps whose coordinates are polar, rotation and center cropping should be applied in the polar coordinates, as shown in Fig. 3(f) and (g) respectively.
@@ 4.4.2 Radar-specific augmentations
In addition to the repurposed subset of vision augmentations, we introduce and experiment with a new domain-specific augmentation for radar SSL we call Radar MIMO Mask (RMM). We briefly explain how the raw data is processed before RMM is applied.
RMM implementation. Several radar formats typically appear in related work: range-azimuth heatmaps, point clouds, or range-Doppler maps [58, 62, 43]. Differently, Radical uses an intermediate 3-D tensor in order to apply RMM augmentations. Specifically, consider a MIMO radar with transmitters and receivers. A range-azimuth heatmap is generated when a preceding 3-D complex tensor is integrated noncoherently over all the antenna pairs (at the first index). RMM is applied before this integration. RMM is best presented as the composition of two operations: (1) antenna dropout, and (2) random phase noise. We further explain these below.
(1) Antenna Dropout. We leverage the reconfigurability of the virtual array emulated by the MIMO radar to design this radar-specific augmentation. We omit randomly a subset of virtual antenna elements from subsequent signal aggregation. Mathematically, we can write
where is the augmented radar heatmap as a function of range and azimuth angle , indexes the set of antenna pairs, and are independent discrete random masks that nullifies the antenna pair with probability . This augmentation simulates scenarios with partial sensor failure or obstructions, which promotes learning from incomplete data and improves robustness. The probability of antenna dropout is a tunable hyper-parameter.
@@
where is the signal from the transmitter and receiver pair, is the augmented signal, and are i.i.d. phase shifts drawn from uniform distributions (in radians), where is a tunable hyper-parameter. This randomization mimics the phase variability introduced by environmental factors and relative motions between the radar and the scene, which is also referred to as Doppler-induced phase noise [43, 28]. Thus it enhances the training coverage of RF conditions likely to occur in the real-world. We note that larger corresponds to more aggressive movements and noise in the environment.
RMM instantiation. The final RMM augmentation is the (order-invariant) composition of the two operations detailed above. We found empirically that the hyper-parameters and lead to the best performance in our experiments (see Sec. 6.2).
4.5 Downstream fine-tuning
@@ 4.6 Implementation details
For the radar backbone, we use Radatron [43], which adopts an FPN-based architecture. The backbone has a two-stream architecture, which takes as inputs high- and low-resolution radar heatmaps. Specifically, each stream goes through a stem layer and then two ResNet stages, which are identical to the building blocks of ResNet50 [30]. Then the two streams are concatenated and fused in a convolutional layer. The resultant feature maps are further encoded via additional ResNet stages, and combined to create the features similar to Detectron2 [70]. We pre-train the backbone of the model (without the FPN and the linear regression heads) as the radar feature extractor. Future research could benefit from changing the backbone[24, 57] and detector[17, 37, 56] architectures.
The vision branch uses a pre-trained CLIP image encoder model [59], which we freeze throughout pre-training.
@@ 5.1 Dataset
We evaluate Radical on the Radatron[43] dataset, which supports raw radar data format. This is because our domain-specific augmentations require raw radar format. In addition to the requisite raw radar format, we find Radatron’s size beneficial in the characterisation we present herein. Out of the unlabeled set, we use K frames for self-supervised pretraining, K annotated frames for supervised fine-tuning, and K annotated frames for testing. The train and test splits are constructed from experiments conducted on different days throughout the data collection campaign. The raw radar frames are first converted to complex, 86-channel heatmaps. These heatmaps are then fed to the network for preprocessing and stochastic augmentation.
@@ 5.2 Experiments
We pre-train Radical as depicted in Fig. 2 and detailed in Secs. 4. We utilize unlabeled radar-vision frames from Radatron as described in Sec. 5.1. We specialize the pre-trained radar embeddings for a downstream task relevant to self-driving. The task strives to detect rotated 2D bounding boxes in BEV from radar heatmaps.
During pre-training, we use a batch size of 64, learning rate of 0.05 and cosine learning rate scheduling with an SGD optimizer with momentum 0.9 and weight decay 0.0001. During fine-tuning, we adopt the same training setting as Radatron, using a batch size of 8, an SGD optimizer with learning rate of 0.01 and 25K iterations with learning rate drop at 15K and 20K iterations. We increase the weight decay to 0.001 in order to avoid overfitting problems, boosting the baseline performance.
Unless otherwise stated, our results are obtained with pre-training the backbone using 32K unlabeled frames, and fine-tuning the downstream model on 13K labeled frames. Results are averaged over 6 runs.
@@ 5.3 Baselines
(2) Intra-modal CL. We disable vision from contributing to the composite contrastive loss, which results in intra-modal, radar-only CL. For this, we use the vision-based augmentations of vertical flipping and center cropping.
@@
Method | |||
Radatron [43] | |||
Intra-modal | |||
Cross-modal | |||
Radical (ours) |
@@ 6 Results
This section presents a comprehensive analysis of Radical’s performance against baselines and examines the impact of various augmentations on model performance.
Evaluation metrics. Following previous radar detection work [58, 43], we use Average Precision (AP) with IoU thresholds of , and to evaluate Radical’s detection performance. We also use the mean AP (mAP) of IoU values from 0.5 to 0.95 in 0.05 steps. We follow the COCO framework [41] for evaluating Radical.
@@
Fine-tuning backbone. We pre-train Radical utilising our composite intra- and cross-modal CL configuration, along with the two baseline CL configurations from Sec. 5.3. We then fine-tune all pre-trained backbones along with their bounding box estimation heads. We compare these pre-trained variants to the implementation of [43] which uses random initialization. Table 1 shows the quantitative results using three metrics: mAP, AP, and AP. The mean and standard deviation of these results are obtained from 6 different runs of supervised training, while keeping the pre-trained weights the same. We see that Radical’s composite intra- and cross-modal configuration performs most favourably, and outperforms random initialization by 5.8% in mAP. This demonstrates the efficacy of Radical’s pre-training on this highly relevant downstream task. Radical also outperforms intra-modal CL and cross-modal CL baselines by 2.9% and 2.4% respectively. Despite good gains over random initialisation (approx. 3% each), the two CL baselines are unable to approach the performance of Radical’s composite CL loss.
Freezing backbone. We freeze the pre-trained weights in order to assess and compare the quality of the learnt features across our contrastive configurations. To this end, we train task-specific heads for our downstream bounding box estimation task similar to above. For Radatron, we randomly initialize its backbone and then similarly freeze it. The averages and standard deviations are listed in Table 2. We observe that Radical outperforms all baselines on all metrics. Random initialisation performs poorly compared to the pre-trained variants. This highlights the inadequacy of the task-specific head to perform accurate bounding box estimation without quality featurisation underneath. We also observe that the gap between Radical and the two CL baselines widens compared to Table 1. Without fine-tuning to compensate, this further underscores the efficacy of Radical compared to the CL baselines. We also observe a slight performance advantage to cross-modal over the intra-modal CL. This could point to the importance of visual priors in the training of quality radar embeddings.
Labeled data fraction | |||
Radatron [43] | |||
Radical (ours) | |||
@@
Method | |||
RMM | |||
Rotation | |||
Center Cropping | |||
Horizontal Flip | |||
Cross-modal (No Aug.) | |||
Threshold | |||
Cutout | |||
Vertical Flip |
Label efficiency. We investigate the impact of the number of available labeled data on performance under a fine-tuning protocol. We compare Radical to the random initialization of Radatron. Table 3 shows the mAP after full fine-tuning as a function of the fraction of labeled data used. We see increasing improvements using Radical pre-training over the supervised baseline across decreased label density regimes.